if i have 12' abs pipe want cut 3 pices one at 33 3/8 and seccon at 56 5/8 and tird at 39 7/8 the saw at 1/8 what got left

Answer by Alan3354(69443)

if i have 12' abs pipe want cut 3 pices one at 33 3/8 and seccon at 56 5/8 and tird at 39 7/8 the saw at 1/8 what got left
------------
12' = 144 inches
----
33 3/8 + 56 5/8 + 39 7/8 = 112 15/8
= 113" 7/8
2 cuts = 1/4"
--> 114" 1/8
-------------
144" - 114 7/8 = 29 1/8